# Segment Any 3D Gaussians

### v2

@@ Abstract
This paper presents SAGA (Segment Any 3D GAussians), a highly efficient 3D promptable segmentation method based on 3D Gaussian Splatting (3D-GS). Given 2D visual prompts as input, SAGA can segment the corresponding 3D target represented by 3D Gaussians within 4 ms. This is achieved by attaching an scale-gated affinity feature to each 3D Gaussian to endow it a new property towards multi-granularity segmentation. Specifically, a scale-aware contrastive training strategy is proposed for the scale-gated affinity feature learning. It 1) distills the segmentation capability of the Segment Anything Model (SAM) from 2D masks into the affinity features and 2) employs a soft scale gate mechanism to deal with multi-granularity ambiguity in 3D segmentation through adjusting the magnitude of each feature channel according to a specified 3D physical scale. Evaluations demonstrate that SAGA achieves real-time multi-granularity segmentation with quality comparable to state-of-the-art methods. As one of the first methods addressing promptable segmentation in 3D-GS, the simplicity and effectiveness of SAGA pave the way for future advancements in this field. Our code will be released.
${}^{\textrm{\Letter}}$orresponding author: wei.shen@sjtu.edu.cn
@@ 1 Introduction
Promptable segmentation has attracted increasing attention and has seen significant advancements, particularly with the development of 2D segmentation foundation models such as the Segment Anything Model (SAM) [24]. However, 3D promptable segmentation remains relatively unexplored due to the scarcity of 3D data and the high cost of annotation. To address these challenges, many studies [7, 10, 53, 23, 11, 35] have proposed to extend SAM’s 2D segmentation capabilities to 3D using radiance fields, achieving notable success.
In this paper, we focus on promptable segmentation in 3D Gaussian Splatting (3D-GS) [21], which represents a significant milestone in radiance fields research due to its superior rendering quality and efficiency compared to its predecessors. We highlight that, in contrast to previous radiance fields, the explicit 3D Gaussian structure can be an ideal carrier for 3D segmentation, as segmentation capabilities can be integrated into 3D-GS as an intrinsic attribute, without necessitating an additional bulky segmentation module.
@@
To address the two challenges, SAGA respectively introduces two solutions. First, SAGA attaches an affinity feature to each 3D Gaussian in a scene to endow it with a new property towards segmentation. The similarity between two affinity features indicates whether the corresponding 3D Gaussians belong to the same 3D target. Second, inspired by GARField [23], SAGA employs a soft scale gate mechanism to handle multi-granularity ambiguity. Depending on a specified 3D physical scale, the scale gate adjusts the magnitude of each feature channel. This mechanism maps the Gaussian affinity features into different sub-spaces for various scales, thereby preserving the multi-granularity information and meanwhile mitigating the distraction in feature learning brought by multi-granularity ambiguity. More specifically, to realize the two solutions, SAGA proposes a scale-aware contrastive training strategy, which distills the segmentation capability of SAM from 2D masks into scale-gated affinity features. This strategy determines the correlations between each pair of pixels within an image based on given 3D scales. These correlations are then used to supervise the rendered affinity features through a correspondence distillation loss. The correlation information is transmitted to the Gaussian affinity features via backpropagation facilitated by the differentiable rasterization algorithm. After training, SAGA achieves real-time multi-granularity segmentation precisely.
Figure 1: SAGA performs promptable multi-granularity segmentation within milliseconds. Prompts are points marked by colors.
Thanks to the native integration of 3D-GS, SAGA is quite efficient. As illustrated in Figure 1, by taking 2D prompts as input, SAGA can segment each 3D target within 4 ms, making it the fastest method currently available. Evaluations on existing benchmarks demonstrate that the segmentation quality of SAGA is on par with previous state-of-the-art methods. Furthermore, with well-trained affinity features, SAGA is capable of automatic scene decomposition and open-vocabulary segmentation. We believe the simplicity and effectiveness of SAGA can pave the way for future advancements in promptable segmentation within 3D-GS.
@@ 2 Related Work
The task of 2D promptable segmentation is proposed by Kirillov et al. [24]. The goal of this task is to return segmentation masks given input prompts that specify the segmentation target in an image. To address this problem, they introduce the Segment Anything Model (SAM), a groundbreaking segmentation foundation model. A similar model to SAM is SEEM [56], which also achieves competitive performance. Prior to these models, the most closely related task to promptable 2D segmentation is interactive image segmentation, which has been explored in numerous studies [4, 16, 17, 44, 9, 45, 31].
@@
Inspired by the success of radiance fields [37, 48, 8, 1, 39, 19, 13, 51, 29, 12], numerous studies have explored 3D segmentation within them. Zhi et al. [55] proposed Semantic-NeRF, demonstrating the potential of Neural Radiance Field (NeRF) in semantic propagation and refinement. NVOS [43] introduced an interactive approach to select 3D objects from NeRF by training a lightweight multi-layer perceptron (MLP) using custom-designed 3D features. By using 2D self-supervised models, approaches like N3F [49], DFF [26], and ISRF [15] aim to elevate 2D visual features to 3D by training additional feature fields that can output 2D feature maps, imitating the original 2D features from different views. NeRF-SOS [11] distills the 2D feature similarities into 3D features. There are also some other instance segmentation and semantic segmentation approaches [46, 40, 54, 33, 34, 3, 14, 50] for radiance fields. Combined with CLIP, some approaches [22, 30, 2, 41] proposed to conduct open-vocabulary 3D segmentation in radiance fields. With the popularity of SAM, a stream of studies [23, 53, 52, 7] proposed lifting the segmentation ability of SAM to 3D with radiance fields. SA3D [7] adopts an iterative pipeline to refine the 3D mask grids with SAM. GaussianGrouping [52] uses video tracking technology to align the inconsistent 2D masks extracted by SAM across different views and assigns labels to 3D Gaussians in a 3D-GS model with the aligned masks. OmniSeg3D [53] employs a hierarchical contrastive learning method to automatically learn hierarchical segmentation from multi-view 2D masks extracted by SAM.
The approach most closely related to SAGA is GARField [23], which uses 3D physical scale as a condition for 3D feature fields to address the multi-granularity ambiguity in 3D segmentation. This concept inspires the scale gate mechanism of SAGA. However, GARField relies on an implicit feature field to output features in 3D space. For each segmentation at different scales, GARField must repeatedly query the scale-conditioned field, significantly hindering segmentation efficiency. In contrast, the scale-gate mechanism of SAGA can be applied to 3D-GS without introducing additional neural networks, thereby enhancing SAGA’s efficiency.
@@ 3 Method
In this section, we first give a brief review of 3D Gaussian Splatting (3D-GS) [21] and the scale-conditioned 3D features [22, 23]. Then we introduce the overall pipeline of SAGA, followed by explanation of the scale-gated Gaussian affinity features and the scale-aware contrastive learning.
@@ 3.1 Preliminary
LERF [22] first proposes the concept of a scale-conditioned feature field for learning from global image embeddings obtained from CLIP. GARField [23] then introduces it into the area of radiance field segmentation to tackle the multi-granularity ambiguity. To compute the 3D mask scale $s_{\mathbf{M}}$ of a 2D mask $\mathbf{M}$, GARField projects $\mathbf{M}$ into 3D space with the camera intrinsic parameters and depth information predicted by a pre-trained radiance field. Let $\mathcal{P}$ denote the obtained point cloud, $\mathcal{X}(\mathcal{P}),\mathcal{Y}(\mathcal{P}),\mathcal{Z}(\mathcal{P})$ denote the set of 3D coordinate components of $\mathcal{P}$, the mask scale $s_{\mathbf{M}}$ is:
@@ 3.2 Overall Pipeline
To train the affinity features, SAGA extracts a set of multi-granularity masks $\mathcal{M}_{\mathbf{I}}=\{\mathbf{M}_{\mathbf{I}}^{i}\in\{0,1\}^{HW}|i=1,...,N_{\mathbf{I}}\}$ for each image $\mathbf{I}$ in the training set $\mathcal{I}$ with SAM. $H,W$ denotes the height and width of $\mathbf{I}$ respectively. $N_{\mathbf{I}}$ denotes the number of extracted masks. For each mask $\mathbf{M}_{\mathbf{I}}^{i}$, its corresponding 3D physical scale $s_{\mathbf{M}_{\mathbf{I}}^{i}}$ is calculated using the depth predicted by $\mathcal{G}$ with the camera pose, as shown in Equation (2). Subsequently, SAGA employs a scale-aware contrastive learning strategy (Section 3.3.2) to distill the multi-granularity segmentation ability embedded in multi-view 2D masks into the scale-gated affinity features. After training, at given scales, the affinity feature similarities between two Gaussians indicate whether they belong to the same 3D target.
During inference (Section 3.4), given a specific viewpoint, SAGA takes 2D visual prompts (with scales) as input and converts them to corresponding 3D scale-gated query features, which are then used to segment the 3D target by evaluating feature similarites with scale-gated 3D affinity features. Additionally, with the well-trained affinity features, 3D scene decomposition can be achieved through straightforward clustering. Furthermore, by integrating with CLIP, SAGA can perform open-vocabulary segmentation (Section A.3) without the need for additional language fields.
3.3 Gaussian Affinity Feature Learning
At the core of SAGA is the Gaussian affinity features =$\mathcal{F}=\{\mathbf{f}_{\mathbf{g}}\mid\mathbf{g}\in\mathcal{G}\}$, which are learned from the multi-view 2D masks extracted by SAM. To tackle the inherent multi-granularity ambiguity in promptable segmentation, inspired by GARfield [23], we introduce a scale gate mechanism to split the feature space into different sub-spaces for various 3D physical scales. Then, a 3D Gaussian can belong to different segmentation targets at different granularities without conflict. To get these scale-gated features, we adopt a scale-aware contrastive learning strategy.
3.3.1 Scale-Gated Affinity Features
Given a Gaussian affinity feature $\mathbf{f}_{\mathbf{g}}$ and a specific scale $s$, SAGA adopts a scale gate to adapt the magnitude of different feature channels accordingly. The scale gate is defined as a mapping $\mathcal{S}:[0,1]\rightarrow[0,1]^{D}$, which projects a scale scalar $s\in[0,1]$ to its corresponding soft gate vector $\mathcal{S}(s)$. To maximize segmentation efficiency, the scale gate adopts a extremely streamlined design, which is composed with a single linear layer followed by a sigmoid function. At the scale of $s$ the scale-gated affinity feature is:
@@ 3.3.1 Scale-Gated Affinity Features
where $\odot$ denotes the Hadamard product. Thanks to the simplicity of the scale gate mechanism, the time overhead caused by scale changing is negligible.
@@
In practice, we find that there are many noisy Gaussians in the 3D space that exhibit unexpectedly high feature similarities with the segmentation target. This may occur for various reasons, such as insufficient training due to small weights in rasterization or incorrect geometry structure learned by 3D-GS. To tackle this problem, we adopt the spatial locality prior of 3D Gaussians. During training, SAGA uses the smoothed affinity feature of a Gaussian $\mathbf{g}$ to replace its original feature $\mathbf{f}_{\mathbf{g}}$, i.e., $\mathbf{f}_{\mathbf{g}}\leftarrow\frac{1}{K}\sum_{\mathbf{g}^{\prime}\in\texttt{KNN}(\mathbf{g})}\mathbf{f}_{\mathbf{g}^{\prime}}$. $\texttt{KNN}(\mathbf{g})$ denotes K-nearest neighbors of $\mathbf{g}$.
After training, the affinity feature for each 3D Gaussian is saved as its smoothed feature. Consequently, there is no need to perform KNN during inference, thus avoiding additional computational overhead.
3.3.2 Scale-Aware Contrastive Learning
To train the scale-gated affinity features, for an image $\mathbf{I}$ of a specific view, we first convert the automatically extracted 2D masks $\mathcal{M}_{\mathbf{I}}$ to scale-aware supervision signal.
@@ 3.3.2 Scale-Aware Contrastive Learning
We assign a scale-aware pixel identity vector $\mathbf{V}(s,\mathbf{p})\in\{0,1\}^{N_{\mathbf{I}}}$ to each pixel $\mathbf{p}$ in an image $\mathbf{I}$. The identity vectors reflect the 2D masks that a pixel belong to at specific scales. Intuitively, if two pixels $\mathbf{p}_{1},\mathbf{p}_{2}$ share at least a same mask at a given scale (i.e., $\mathbf{V}(s,\mathbf{p}_{1})\cdot\mathbf{V}(s,\mathbf{p}_{2})>0$), they should have similar features at scale $s$.
To obtain $\mathbf{V}(s,\mathbf{p})$, we first sort the mask set $\mathcal{M}_{\mathbf{I}}$ in descending order according to their mask scales and get an ordered mask list $\mathcal{O}_{\mathbf{I}}=(\mathbf{M}_{\mathbf{I}}^{(1)},...,\mathbf{M}_{\mathbf{I}}^{(N_{\mathbf{I}})})$, where $s_{\mathbf{M}_{\mathbf{I}}^{(1)}}>...>s_{\mathbf{M}_{\mathbf{I}}^{(N_{\mathbf{I}})}}$. Then, for a pixel $\mathbf{p}$, when $s_{\mathbf{M}^{(i)}_{\mathbf{I}}}<s$, the i-th entry of $\mathbf{V}(s,\mathbf{p})$ is set to $\mathbf{M}^{(i)}_{\mathbf{I}}(\mathbf{p})$. When $s_{\mathbf{M}^{(i)}_{\mathbf{I}}}\geq s$, the i-th entry of $\mathbf{V}(s,\mathbf{p})$ equals to $1$ only if $\mathbf{M}^{(i)}_{\mathbf{I}}(\mathbf{p})=1$ and all smaller masks in $\{\mathbf{M}^{(j)}_{\mathbf{I}}\mid s\leq\mathbf{M}^{(j)}_{\mathbf{I}}<s_{\mathbf{M}^{(i)}_{\mathbf{I}}}\}$ equals to 00 at pixel $\mathbf{p}$. Formally, we have:
$\mathbf{V}^{i}(s,\mathbf{p})=\left\{\begin{array}[]{cl}\mathbf{M}^{(i)}_{\mathbf{I}}(\mathbf{p})&\text{if }s_{\mathbf{M}^{(i)}_{\mathbf{I}}}<s\text{ or}\left(\forall\ \mathbf{M}\in\{\mathbf{M}^{(j)}_{\mathbf{I}}\mid s\leq s_{\mathbf{M}^{(j)}_{\mathbf{I}}}<s_{\mathbf{M}^{(i)}_{\mathbf{I}}}\},\mathbf{M}(\mathbf{p})=0\right)\\ 0&\text{otherwise}\end{array}\right.$ (6)
This assignment of pixel identity vectors is based on the assumption that if a pixel belongs to a specific mask at a given scale, it will continue to belong to that mask at larger scales.
@@
We adapt the correspondence distillation loss proposed by Hamilton et al. [18] for training the scale-gated Gaussian affinity features. Concretely, for two pixels $\mathbf{p}_{1},\mathbf{p}_{2}$ at a given scale $s$, their mask correspondence is given by:
$\texttt{Corr}_{m}(s,\mathbf{p}_{1},\mathbf{p}_{2})=\mathbbm{1}(\mathbf{V}(s,\mathbf{p}_{1})\cdot\mathbf{V}(s,\mathbf{p}_{2})),$ (7)
where $\mathbbm{1}(\cdot)$ is the indicator function, which is equal to 1 when the input greater than 0 and equal to 0 when the input is 0. The feature correspondence between two pixels is defined as the cosine similarity between their scale-gated features:
@@
The correspondence distillation loss $\mathcal{L}_{\texttt{corr}}(\mathbf{p}_{1},\mathbf{p}_{2})$ between two pixels is given by [^1] The feature correspondence is clipped at 0 to stabilize training. Please refer to [18] for more details.:
$\mathcal{L}_{\texttt{corr}}(s,\mathbf{p}_{1},\mathbf{p}_{2})=(1-2\ \texttt{Corr}_{m}(s,\mathbf{p}_{1},\mathbf{p}_{2}))\max(\texttt{Corr}_{f}(s,\mathbf{p}_{1},\mathbf{p}_{2}),0).$ (9)
@@
$\mathcal{L}=\sum_{(\mathbf{p}_{1},\mathbf{p}_{2})\in\delta(\mathbf{I})\times\delta(\mathbf{I})}\mathcal{L}_{\texttt{corr}}(\mathbf{p}_{1},\mathbf{p}_{2})+\frac{1}{HW}\sum_{\mathbf{p}\in\delta(\mathbf{I})}\mathcal{L}_{\texttt{norm}}(\mathbf{p}),$ (12)
@@
3.3.3 Additional Training Strategy
During the training of SAGA, an unavoidable issue is the data imbalance, which is reflected in three aspects: 1) Most pixel pairs keep to be positive or negative regardless of scale variations, making the learned feature insensitive to scales; 2) The majority of pixel pairs shows negative correspondence, resulting in feature collapse; 3) Large targets that occupy more pixels in images have more effect on the optimization, leading to bad performance of segmenting small targets. We tackle this problem by resampling the sampled pixel-pairs and re-weighting the loss function for different samples. Please refer to Section A.1 for detailed explanation.

### v3

@@ Abstract
This paper presents SAGA (Segment Any 3D GAussians), a highly efficient 3D promptable segmentation method based on 3D Gaussian Splatting (3D-GS). Given 2D visual prompts as input, SAGA can segment the corresponding 3D target represented by 3D Gaussians within 4 ms. This is achieved by attaching an scale-gated affinity feature to each 3D Gaussian to endow it a new property towards multi-granularity segmentation. Specifically, a scale-aware contrastive training strategy is proposed for the scale-gated affinity feature learning. It 1) distills the segmentation capability of the Segment Anything Model (SAM) from 2D masks into the affinity features and 2) employs a soft scale-gate mechanism to deal with multi-granularity ambiguity in 3D segmentation through adjusting the magnitude of each feature channel according to a specified 3D physical scale. Evaluations demonstrate that SAGA achieves real-time multi-granularity segmentation with quality comparable to state-of-the-art methods. As one of the first methods addressing promptable segmentation in 3D-GS, the simplicity and effectiveness of SAGA pave the way for future advancements in this field.
Codehttps://github.com/Jumpat/SegAnyGAussians
Figure 1: SAGA performs promptable multi-granularity segmentation within milliseconds. Prompts are marked by points.
@@ 1 Introduction
Promptable segmentation has attracted increasing attention and has seen significant advancements, particularly with the development of 2D segmentation foundation models such as the Segment Anything Model (SAM) (Kirillov et al. 2023). However, 3D promptable segmentation remains relatively unexplored due to the scarcity of 3D data and the high cost of annotation. To address these challenges, many studies (Cen et al. 2023; Chen et al. 2023; Ying et al. 2024; Kim et al. 2024; Fan et al. 2023; Liu et al. 2024b) have proposed to extend SAM’s 2D segmentation capabilities to 3D using radiance fields, achieving notable success.
In this paper, we focus on promptable segmentation in 3D Gaussian Splatting (3D-GS) (Kerbl et al. 2023), which represents a significant milestone in radiance fields research due to its superior rendering quality and efficiency compared to its predecessors. We highlight that, in contrast to previous radiance fields, the explicit 3D Gaussian structure is an ideal carrier for 3D segmentation, as segmentation capabilities can be integrated into 3D-GS as an intrinsic attribute, without necessitating an additional bulky segmentation module.
@@
To address the two challenges, SAGA respectively introduces two solutions. First, SAGA attaches an affinity feature to each 3D Gaussian in a scene to endow it with a new property towards segmentation. The similarity between two affinity features indicates whether the corresponding 3D Gaussians belong to the same 3D target. Second, inspired by GARField (Kim et al. 2024), SAGA employs a soft scale-gate mechanism to handle multi-granularity ambiguity. Depending on a specified 3D physical scale, the scale gate adjusts the magnitude of each feature channel. This mechanism maps the Gaussian affinity features into different sub-spaces for various scales, thereby preserving the multi-granularity information and meanwhile mitigating the distraction in feature learning brought by multi-granularity ambiguity. To realize the two solutions, SAGA proposes a scale-aware contrastive training strategy, which distills the segmentation capability of SAM from 2D masks into scale-gated affinity features. This strategy determines the correlations between a pair of pixels within an image based on 3D scales. These correlations are then used to supervise the rendered affinity features through a correspondence distillation loss. The correlation information is transmitted to the Gaussian affinity features via backpropagation facilitated by the differentiable rasterization algorithm. After training, SAGA achieves real-time multi-granularity segmentation precisely.
@@ 2 Related Work
The task of 2D promptable segmentation is proposed by Kirillov et al. (2023), which aims to return segmentation masks given input prompts that specify the segmentation target in an image. To address this problem, they introduce the Segment Anything Model (SAM), a groundbreaking segmentation foundation model. A similar model to SAM is SEEM (Zou et al. 2023), which also achieves competitive performance. Prior to these models, the most closely related task to promptable 2D segmentation is interactive image segmentation (Boykov and Jolly 2001; Grady 2006; Gulshan et al. 2010; Rother, Kolmogorov, and Blake 2004; Chen et al. 2022b; Sofiiuk, Petrov, and Konushin 2022; Liu et al. 2023b). Inspired by the success of SAM, many studies (Yang et al. 2023; Xu et al. 2023; Guo et al. 2024; Yin et al. 2024) proposed to use SAM for 3D segmentation. Different from them, we focus on lifting the ability of SAM to 3D via 3D-GS.
@@
With the success of radiance fields (Mildenhall et al. 2020; Sun, Sun, and Chen 2022; Chen et al. 2022a; Barron et al. 2022; Müller et al. 2022; Hedman et al. 2024; Fridovich-Keil et al. 2022; Wizadwongsa et al. 2021; Lindell, Martel, and Wetzstein 2021; Fang et al. 2022), numerous studies have explored 3D segmentation within them. Zhi et al. (2021) proposed Semantic-NeRF, demonstrating the potential of Neural Radiance Field (NeRF) in semantic propagation and refinement. NVOS (Ren et al. 2022) introduced an interactive approach to select 3D objects from NeRF by training a lightweight MLP using custom-designed 3D features. By using 2D self-supervised models, approaches like N3F (Tschernezki et al. 2022), DFF (Kobayashi, Matsumoto, and Sitzmann 2022), and ISRF (Goel et al. 2023) aim to elevate 2D visual features to 3D by training additional feature fields that can output 2D feature maps, imitating the original 2D features from different views. NeRF-SOS (Fan et al. 2023) and ContrastiveLift (Bhalgat et al. 2023) distill the 2D feature similarities into 3D features. There are also some other instance segmentation and semantic segmentation approaches (Stelzner, Kersting, and Kosiorek 2021; Niemeyer and Geiger 2021; Yu, Guibas, and Wu 2022; Liu et al. 2022, 2023c; Bing, Chen, and Yang 2023; Fu et al. 2022; Vora et al. 2022; Siddiqui et al. 2023) for radiance fields. Combined with CLIP (Radford et al. 2021), some approaches (Kerr et al. 2023; Liu et al. 2023a; Bhalgat et al. 2024; Qin et al. 2024) proposed to conduct open-vocabulary 3D segmentation in radiance fields. With the popularity of SAM, a stream of studies (Kim et al. 2024; Ying et al. 2024; Ye et al. 2024; Cen et al. 2023; Lyu et al. 2024) proposed lifting the segmentation ability of SAM to 3D with radiance fields. SA3D (Cen et al. 2023) adopts an iterative pipeline to refine the 3D mask grids with SAM. GaussianGrouping (Ye et al. 2024) uses video tracking technology to align the inconsistent 2D masks extracted by SAM across different views and assigns labels to 3D Gaussians in a 3D-GS model with the aligned masks. OmniSeg3D (Ying et al. 2024) employs a hierarchical contrastive learning method to automatically learn segmentation from multi-view 2D masks extracted by SAM.
The approach most closely related to SAGA is GARField (Kim et al. 2024), which addresses multi-granularity ambiguity in 3D segmentation using 3D physical scale, inspiring SAGA’s scale gate mechanism. However, GARField’s reliance on implicit feature fields for outputting 3D features requires repeated queries for segmentation at different scales, reducing efficiency. In contrast, the scale-gate mechanism of SAGA enhances efficiency by integrating directly with 3D-GS without additional computation.
@@ 3 Method
In this section, we first give a brief review of 3D Gaussian Splatting (3D-GS) (Kerbl et al. 2023) and the scale-conditioned 3D features (Kerr et al. 2023; Kim et al. 2024). Then we introduce the overall pipeline of SAGA, followed by explanation of the scale-gated Gaussian affinity features and the scale-aware contrastive learning.
@@ 3.1 Preliminary
LERF (Kerr et al. 2023) first proposes the concept of a scale-conditioned feature field for learning from global image embeddings obtained from CLIP. GARField (Kim et al. 2024) then introduces it into the area of radiance field segmentation to tackle the multi-granularity ambiguity. To compute the 3D mask scale $s_{\mathbf{M}}$ of a 2D mask $\mathbf{M}$, GARField projects $\mathbf{M}$ into 3D space with the camera intrinsic parameters and depth information predicted by a pre-trained radiance field. Let $\mathcal{P}$ denote the obtained point cloud, $\mathcal{X}(\mathcal{P}),\mathcal{Y}(\mathcal{P}),\mathcal{Z}(\mathcal{P})$ denote the set of 3D coordinate components of $\mathcal{P}$, the mask scale $s_{\mathbf{M}}$ is:
@@ 3.2 Overall Pipeline
To train the affinity features, SAGA extracts a set of multi-granularity masks $\mathcal{M}_{\mathbf{I}}=\{\mathbf{M}_{\mathbf{I}}^{i}\in\{0,1\}^{HW}|i=1,...,N_{\mathbf{I}}\}$ for each image $\mathbf{I}$ in the training set $\mathcal{I}$ with SAM. $H,W$ denotes the height and width of $\mathbf{I}$ respectively. $N_{\mathbf{I}}$ is the number of extracted masks. For each mask $\mathbf{M}_{\mathbf{I}}^{i}$, its 3D physical scale $s_{\mathbf{M}_{\mathbf{I}}^{i}}$ is calculated using the depth predicted by $\mathcal{G}$ with the camera pose, as shown in Equation (2). Subsequently, SAGA employs a scale-aware contrastive learning strategy (Section 3.4) to distill the multi-granularity segmentation ability embedded in multi-view 2D masks into the scale-gated affinity features. After training, at given scales, the affinity feature similarities between two Gaussians indicate whether they belong to the same 3D target.
During inference (Section 3.5), given a specific viewpoint, SAGA converts 2D visual prompts (points with scales) into corresponding 3D scale-gated query features to segment the 3D target by evaluating feature similarities with 3D affinity features. Additionally, with well-trained affinity features, 3D scene decomposition is achievable through simple clustering. Moreover, by integrating with CLIP, SAGA can perform open-vocabulary segmentation (see Section A.3 of the supplement) without requiring language fields.
3.3 Gaussian Affinity Feature
At the core of SAGA is the Gaussian affinity features =$\mathcal{F}=\{\mathbf{f}_{\mathbf{g}}\mid\mathbf{g}\in\mathcal{G}\}$, which are learned from the multi-view 2D masks extracted by SAM. To tackle the inherent multi-granularity ambiguity in promptable segmentation, inspired by GARfield (Kim et al. 2024), we introduce a scale-gate mechanism to split the feature space into different sub-spaces for various 3D physical scales. Then, a 3D Gaussian can belong to different segmentation targets at different granularities without conflict.
Scale-Gated Affinity Features
Given a Gaussian affinity feature $\mathbf{f}_{\mathbf{g}}$ and a specific scale $s$, SAGA adopts a scale gate to adapt the magnitude of different feature channels accordingly. The scale gate is defined as a mapping $\mathcal{S}:[0,1]\rightarrow[0,1]^{D}$, which projects a scale scalar $s\in[0,1]$ to its corresponding soft gate vector $\mathcal{S}(s)$. To maximize segmentation efficiency, the scale gate adopts a extremely streamlined design, which is a single linear layer followed by a sigmoid function. At the scale of $s$ the scale-gated affinity feature is:
@@ 3.3 Gaussian Affinity Feature
where $\odot$ denotes the Hadamard product. Thanks to the simplicity of the scale-gate mechanism, the time overhead caused by scale changing is negligible.
@@
In practice, we find that there are many noisy Gaussians in the 3D space that exhibit unexpectedly high feature similarities with the segmentation target. This may occur for various reasons, such as insufficient training due to small weights in rasterization or incorrect geometry structure learned by 3D-GS. To tackle this problem, we adopt the spatial locality prior of 3D Gaussians. During training, SAGA uses the smoothed affinity feature of a Gaussian $\mathbf{g}$ to replace its original feature $\mathbf{f}_{\mathbf{g}}$, i.e., $\mathbf{f}_{\mathbf{g}}\leftarrow\frac{1}{K}\sum_{\mathbf{g}^{\prime}\in\texttt{KNN}(\mathbf{g})}\mathbf{f}_{\mathbf{g}^{\prime}}$. $\texttt{KNN}(\mathbf{g})$ denotes K-nearest neighbors of $\mathbf{g}$. After training, the affinity feature for each 3D Gaussian is saved as its smoothed feature.
3.4 Scale-Aware Contrastive Learning
As introduced in Section 3.3, each 3D Gaussian $\mathbf{g}$ is assigned with an affinity feature $\mathbf{f}_{\mathbf{g}}$. To train these features, we employ a scale-aware contrastive learning strategy to distill the pixel-wise correlation information from 2D masks into 3D Gaussians via the differentiable rasterization.
@@ 3.4 Scale-Aware Contrastive Learning
To conduct scale-aware contrastive learning, for an image $\mathbf{I}$, we first convert the automatically extracted 2D masks $\mathcal{M}_{\mathbf{I}}$ to scale-aware supervision signal. For this purpose, we assign a scale-aware pixel identity vector $\mathbf{V}(s,\mathbf{p})\in\{0,1\}^{N_{\mathbf{I}}}$ to each pixel $\mathbf{p}$ in $\mathbf{I}$. The identity vectors reflect the 2D masks that a pixel belong to at specific scales. If two pixels $\mathbf{p}_{1},\mathbf{p}_{2}$ share at least a same mask at a given scale (i.e., $\mathbf{V}(s,\mathbf{p}_{1})\cdot\mathbf{V}(s,\mathbf{p}_{2})>0$), they should have similar features at scale $s$.
To obtain $\mathbf{V}(s,\mathbf{p})$, we first sort the mask set $\mathcal{M}_{\mathbf{I}}$ in descending order according to their mask scales and get an ordered mask list $\mathcal{O}_{\mathbf{I}}=(\mathbf{M}_{\mathbf{I}}^{(1)},...,\mathbf{M}_{\mathbf{I}}^{(N_{\mathbf{I}})})$, where $s_{\mathbf{M}_{\mathbf{I}}^{(1)}}>...>s_{\mathbf{M}_{\mathbf{I}}^{(N_{\mathbf{I}})}}$. Then, for a pixel $\mathbf{p}$, when $s_{\mathbf{M}^{(i)}_{\mathbf{I}}}<s$, the i-th entry of $\mathbf{V}(s,\mathbf{p})$ is set to $\mathbf{M}^{(i)}_{\mathbf{I}}(\mathbf{p})$. When $s_{\mathbf{M}^{(i)}_{\mathbf{I}}}\geq s$, the i-th entry of $\mathbf{V}(s,\mathbf{p})$ equals to $1$ only if $\mathbf{M}^{(i)}_{\mathbf{I}}(\mathbf{p})=1$ and all smaller masks in $\{\mathbf{M}^{(j)}_{\mathbf{I}}\mid s\leq s_{\mathbf{M}^{(j)}_{\mathbf{I}}}<s_{\mathbf{M}^{(i)}_{\mathbf{I}}}\}$ equals to 00 at pixel $\mathbf{p}$. Formally, we have:
$\mathbf{V}^{i}(s,\mathbf{p})=\begin{cases}\mathbf{M}^{(i)}_{\mathbf{I}}(\mathbf{p})&\text{if }s_{\mathbf{M}^{(i)}_{\mathbf{I}}}<s\text{ or}\mathcal{C}(\mathbf{p})\\ 0&\text{otherwise}\end{cases}$ (6)
$\mathcal{C}(\mathbf{p})\triangleq\left(\forall\mathbf{M}\in\{\mathbf{M}^{(j)}_{\mathbf{I}}\mid s\leq s_{\mathbf{M}^{(j)}_{\mathbf{I}}}<s_{\mathbf{M}^{(i)}_{\mathbf{I}}}\},\mathbf{M}(\mathbf{p})=0\right)$
This assignment of pixel identity vectors is based on the fact that if a pixel belongs to a specific mask at a given scale, it will continue to belong to that mask at larger scales.
@@
We adapt the correspondence distillation loss (Hamilton et al. 2022) for training the scale-gated Gaussian affinity features. Concretely, for two pixels $\mathbf{p}_{1},\mathbf{p}_{2}$ at a given scale $s$, their mask correspondence is given by:
$\texttt{Corr}_{m}(s,\mathbf{p}_{1},\mathbf{p}_{2})=\mathds{1}(\mathbf{V}(s,\mathbf{p}_{1})\cdot\mathbf{V}(s,\mathbf{p}_{2})),$ (7)
where $\mathds{1}(\cdot)$ is the indicator function, which is equal to 1 when the input greater than or equal to 0. The feature correspondence between two pixels is defined as the cosine similarity between their scale-gated features:
@@
The correspondence distillation loss $\mathcal{L}_{\texttt{corr}}(\mathbf{p}_{1},\mathbf{p}_{2})$ between two pixels is given by [^1] The feature correspondence is clipped at 0 to stabilize training. Please refer to Hamilton et al. (2022) for more details.:
$\mathcal{L}_{\texttt{corr}}(s,\mathbf{p}_{1},\mathbf{p}_{2})=$ $\ (1-2\cdot\texttt{Corr}_{m}(s,\mathbf{p}_{1},\mathbf{p}_{2}))$ (9)
$\cdot\max(\texttt{Corr}_{f}(s,\mathbf{p}_{1},\mathbf{p}_{2}),0)$
@@
$\mathcal{L}=\sum_{(\mathbf{p}_{1},\mathbf{p}_{2})\in\delta(\mathbf{I})\times\delta(\mathbf{I})}\mathcal{L}_{\texttt{corr}}(\mathbf{p}_{1},\mathbf{p}_{2})+\sum_{\mathbf{p}\in\delta(\mathbf{I})}\mathcal{L}_{\texttt{norm}}(\mathbf{p}),$ (12)
@@
Additional Training Strategy
During the training, an unavoidable issue is the data imbalance, reflected by: 1) Most pixel pairs keep to be positive or negative regardless of scale variations, making the learned feature insensitive to scales; 2) The majority of pixel pairs shows negative correspondence, resulting in feature collapse; 3) Large targets that occupy more pixels in images have more effect on the optimization, leading to bad performance of segmenting small targets. We tackle this problem by resampling the pixel-pairs and re-weighting the loss function for different samples. Please refer to Section A.1 of the appendix.